Project S requires an initial outlay at t = 0 of $13,000, and its expected cash flows would be $6,000 per year for 5 years. Mutually exclusive Project L requires an initial outlay at t = 0 of $33,000, and its expected cash flows would be $8,750 per year for 5 years. If both projects have a WACC of 12%, which project would you recommend?

Select the correct answer.

a. Project S, since the NPVS > NPVL.
b. Project L, since the NPVL > NPVS.
c. Both Projects S and L, since both projects have NPV's > 0.
d. Neither Project S nor L, since each project's NPV < 0.
e. Both Projects S and L, since both projects have IRR's > 0.
Answer #1

S:

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=6000[1-(1.12)^-5]/0.12

=6000*3.6047762

=21628.66

NPV=Present value of inflows-Present value of outflows

=21628.66-13000

=$8628.66(Approx).

L:

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=8750[1-(1.12)^-5]/0.12

=8750*3.6047762

=31541.79

NPV=Present value of inflows-Present value of outflows

=31541.79-33000

=$(1458.21)(Approx).(Negative).

Hence S must be selected having higher and positive NPV

Hence the correct option is:

a. Project S, since the NPVS > NPVL